OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

Point Slope form: \[y - y1 = m(x - x1)\]Where y1 and x1 stand for the given x and y coordinates, and m is the slope. So substitute the values in: \[y - (-6) = 4(x - 3)\]\[y + 6 = 4(x - 3)\]

OpenStudy (anonymous):

\[y + 6 = 4(x - 3)\]You want it in Slope-Intercept form: \[y = mx + b\]So lets solve for y. Distribute the 4.\[y + 6 = 4x - 12\]Subtract 6 from both sides. \[y + 6 = 4x - 12\] -6 -6 Now you have:\[y = 4x - 18\]

OpenStudy (anonymous):

Standard form is Ax + By = C where A can NEVER be a fraction, decimal, or a negative number.

OpenStudy (anonymous):

So you have: y = 4x - 18 Subtract 4x from both sides. -4x + y = -18 ^No negatives allowed. So we multiply everything by -1. \[-1(-4x + y = -18)\]-1  So the standard form is: 4x - y = 18